# Yuchen Zhang

### v2

@@ Abstract
The detection and grounding of multimedia manipulation has emerged as a critical challenge in combating AI-generated disinformation. While existing methods have made progress in recent years, we identify two fundamental limitations in current approaches: (1) Underestimation of MLLM-driven deception risk: prevailing techniques primarily address rule-based text manipulations, yet fail to account for sophisticated misinformation synthesized by multimodal large language models (MLLMs) that can dynamically generate semantically coherent, contextually plausible yet deceptive narratives conditioned on manipulated images; (2) Unrealistic misalignment artifacts: currently focused scenarios rely on artificially misaligned content that lacks semantic coherence, rendering them easily detectable. To address these gaps holistically, we propose a new adversarial pipeline that leverages MLLMs to generate high-risk disinformation. Our approach begins with constructing the MLLM-Driven Synthetic Multimodal (MDSM) dataset, where images are first altered using state-of-the-art editing techniques and then paired with MLLM-generated deceptive texts that maintain semantic consistency with the visual manipulations. Building upon this foundation, we present the Artifact-aware Manipulation Diagnosis via MLLM (AMD) framework featuring two key innovations: Artifact Pre-perception Encoding strategy and Manipulation-Oriented Reasoning, to tame MLLMs for the MDSM problem. Comprehensive experiments validate our framework’s superior generalization capabilities as a unified architecture for detecting MLLM-powered multimodal deceptions. In cross-domain testing on the MDSM dataset, AMD achieves the best average performance, with 88.18 ACC, 60.25 mAP, and 61.02 mIoU scores.
[…]
figure Comparison between widely-used $\text{DGM}^{4}$ [DGM4_TPAMI] and MDSM (ours). (left) $\text{DGM}^{4}$ typically treats visual context manipulation as two independent procedures as rule-based text editing and image editing. The absence of context integration often results in poorly aligned samples, which can be readily perceived by the public. (right) In a real-world scenario, we usually face well-wrapped content on purpose. To mimic such a situation, we propose a new manipulation paradigm, which explicitly provides the modified image as well as meta info of facial editing to Multimodal Large Language Model (MLLM). Then we harness MLLM to generate contextually consistent, deceptive texts to form the challenging image-text pairs.
@@ 1 Introduction
Generative techniques have become a transformative force in artificial intelligence [wu2024deepseekvl2, peng2025lmmr1, dawn2024_ICLR, fu2025speculative, Florence2_CVPR, zhang2025ctrl], showing remarkable adaptability across various domains and modalities. These advancements, while enriching multimedia content, also pose significant challenges to information security. In the media industry in particular, maliciously fake content manipulated by such models can profoundly mislead audiences [Defen_fake_news_NeurIPS, deepfake_eccv_shao]. The unchecked spread of fake media has already negatively affected political, financial, and other sectors [cantarella2023does_fakenews, fake_news_impact_financial, rocha2021impact_health], gradually becoming a major social issue [Olan2024_fakeNewsImpact].
@@
While various fake news scenarios have been explored, including [luo_newsclippings] investigation of out-of-context social image-text pairs and [DGM4] work on detecting randomly tampered regions or words. Our analysis reveals two critical limitations in existing research: 1) Neglect of emerging risks from MLLMs: Current paradigms predominantly address rule-based text manipulation, overlooking the sophisticated linguistic capabilities of modern multi-modal large language models (MLLMs). MLLM-generated text exhibits superior fluency and contextual coherence, significantly increasing deception potential and public susceptibility. 2) Semantic misaligned artifacts. Most methodologies independently manipulate visual and textual elements, producing semantically discordant multimedia outputs. This misalignment not only renders manufactured disinformation too simplistic to effectively deceive the public, but also fails to replicate real-world adversarial behavior, as sophisticated attackers typically maintain meticulous visual-textual consistency to maximize manipulative impact. Both limitations render the multi-modal disinformation scenarios considered in existing works insufficiently realistic.
@@
To address these weaknesses, we take MLLM into consideration and focus on detecting the semantic-aligned manipulation. We first construct the MLLM-Driven Synthetic Multi-modal (MDSM) dataset, which manipulates image and text in a coordinated fashion using MLLM. For the image manipulation, we consider the typical Face Swap and Face Attribute editing. For text, We innovatively guide MLLM to generate modality-aligned yet misleading fake narratives using image editing metadata. As shown in Fig. [Uncaptioned image] The Coherence Trap: When MLLM-Crafted Narratives Exploit Manipulated Visual Contexts(b), after replacing Donald Trump’s face with Micheál Martin’s, we use the swapped name, Micheál Martin, to guide MLLM in generating text, ensuring that the named entity in the text aligns with the image. Following this strategy, we construct over 441k sample pairs.
@@
The alignment of modalities and the authentic texts from MLLMs pose significant challenges for the detection of manipulated media. First, the strategy of perceiving inconsistencies between images and text through contrastive learning, as employed by prior works [zhang2024asap, DGM4], is ineffective in MDSM where images and text are well-matched already. Merely observing aligned image-text pairs is inadequate for reliable detection. Consequently, external clues and contextual knowledge are essential. Second, existing architectures like ASAP [zhang2024asap] and HAMMER [DGM4], which feature multiple detection and grounding heads, are complex and lack generalizability to unseen media sources. To address these challenges, we propose Artifact-aware Manipulation Diagnosis via MLLM (AMD), which leverages MLLMs’ comprehensive understanding of real-world multimedia and their ability to provide unified textual outputs. And AMD generates detection and grounding results in a coherent, text-based format, offering a more intuitive and generalized solution. In summary, we highlight our contributions of this paper as follows:
@@
  • We propose an Artifact-aware Manipulation Diagnosis framework (AMD) for the MDSM problem that synergizes artifact pre-perception encoding and manipulation-oriented reasoning to effectively adapt MLLMs for precise manipulation analysis.
@@
2 Related Work
Deepfake Detection. The rapid progress of generative models and the surge in synthetic content have accelerated advances in Deepfake detection. Existing work spans unimodal and multimodal approaches. Unimodal, including image-based [ICCV_9710718, CVPR_9157215] and text-based [ACL_NEP_text_unimodal, huang2023faking] approaches, already achieve strong results. With the rise of Multimodal Large Language Models (MLLMs), multimodal Deepfake detection has received increasing attention [liu2024mmfakebench, DGM4, liu2024fkaowl]. Regarding datasets, Shao et al. [DGM4] introduced the pioneering DGM4 benchmark for multimodal manipulation detection and grounding. However, its manipulations are rule-based, leading to semantically fragmented image-text discrepancies that do not accurately reflect real-world misinformation. MMFakeBench [liu2024mmfakebench] recognized this limitation and proposed generating semantically aligned news images using text-to-image models. Yet such semantically matched samples constitute only 30% of its fake subset, and the dataset contains merely 11k samples, limiting its utility for training robust detectors. Existing Deepfake datasets also fail to consider the risk of semantically coherent but misleading text generated by modern MLLMs. On the methodological side, HAMMER [DGM4] integrates contrastive learning to build a detector capable of classifying manipulation types and grounding manipulated regions, but it does not address cross-domain robustness. Beyond conventional multimodal detectors, FKA-Owl [liu2024fkaowl] employs a 7B-scale MLLM with several architectural modifications to enhance generalization. However, it is trained on DGM4, where text manipulations follow fixed editing rules rather than being synthesized by MLLMs, making it unsuitable for detecting more subtle, semantically aligned misinformation produced by modern models. Moreover, FKA-Owl performs only binary real/fake classification without fine-grained manipulation type prediction or grounding, and its large backbone and heavy architectural design result in substantially increased computational cost and slower inference.
Multi-Modal Large Language Model. In recent years, Multi-Modal Large Language Models have emerged as a crucial technology for understanding and reasoning across multiple modalities, particularly text and images. By extending the capabilities of Large Language Models (LLMs) to incorporate visual inputs, these models have demonstrated outstanding performance in tasks such as image captioning and visual question answering. CLIP  [CLIP_Radford2021LearningTV] and ALIGN  [ALIGN_Jia2021ScalingUV] leveraged contrastive learning to align visual and textual representations, enabling efficient zero-shot vision-language understanding. Subsequently, models such as Flamingo  [Flamingo_NEURIPS2022] and BLIP-2  [BLIP2_ICML] have introduced vision-language transformers, integrating pre-trained LLMs with vision encoders to enhance cross-modal reasoning and generative capabilities. More recently, GPT-4V  [GPT4TR] and Florence-2  [Florence2_CVPR] have significantly enhanced the potential of MLLMs in tackling complex multi-modal tasks by leveraging a more efficient framework and larger-scale pre-training data. A key advantage of MLLMs is their acquisition of extensive world knowledge through large-scale pretraining, which substantially strengthens their reasoning abilities in downstream tasks. Such knowledge not only enhances cross-modal understanding but also proves essential for handling challenging problems, including misinformation detection.
@@ 3.1 Multi-Modal Media Source Collection
We use the GoodNews  [GoodNews], VisualNews  [VisualNews], and N24News  [N24News] datasets as the Source News Pool $\mathcal{D}_{O}$, which consists of over 2.1M image-text pairs sourced from various real-world news outlets. Given the significant influence of human-centric news among various forms of multi-modal media, we focus on human-centric data for MDSM. $\mathcal{D}_{O}$ is firstly filtered by detecting faces in images with Dlib  [dlib] and identifying person names in texts with BERT  [devlin2018bert]. Only pairs, $p_{s}=(I_{s},T_{s})$, containing both faces and named entities are used for manipulation. Additionally, we collect the Celebrity Head-shot Dataset $\mathcal{D}_{C}$, which contains about 30k pairs of head-shot images and corresponding names to facilitate the aligned manipulation for Face Swap. Details are provided in the appendix.
@@ 3.2 Multi-Modal Media Manipulation
In the image modality, two main attacks, Face Swap (FS) and Face Attribute (FA), are employed. For the text modality, we utilize advanced MLLM to generate semantic-aligned texts for the images.
Datasets Samples Modality Tasks Training Semantic MLLM
LIAR  [Text_news_dataset_liar] 13K
DFIM-HQ  [DFIM_HQ] 140K
MEIR  [MEIR_ACM18] 139k
MiRAGeNews  [huang-etal-2024-miragenews] 15k
COSMOS  [COSMOS_AAAI] 453k
MMFakeBench  [liu2024mmfakebench] 11k
$\text{DGM}^{4}$  [DGM4] 230k
MDSM (Ours) 441k
@@
$\triangleright$ Face Swap. Face swap is a critical tool for attackers to forge images of public figures and politicians, posing threats to societal security. We use two representative face swap methods, SimSwap  [chen2020simswap] and e4s  [liu2022fine_e4s], to perform such manipulations. We prioritize modifying larger faces to target the primary subject in the image (Fig. 2c). Given a source image $I_{s}$, we randomly choose one of the two methods and replace the largest face $F_{s}$ in $I_{s}$ with a face $F_{c}$ from $\mathcal{D}_{C}$, generating a manipulated face swap sample $I_{fs}$. The bounding box $y_{box}=\{x_{1},y_{1},x_{2},y_{2}\}$ of the swapped face and the name of $F_{c}$ are recorded. To keep the image-text aligned, corresponding processing is also done on the authentic text. We use MLLM to identify the name of $F_{s}$, as shown in Fig. 1, $F_{s}$ is identified as Trump. We then refine the authentic text for FS category by replacing this name with $F_{c}$’s name.
$\triangleright$ Face Swap and Text Fabrication. We use Qwen2-VL  [Qwen2-VL] to generate consistent but misleading narratives. This requires knowing: 1) the inserted person’s name, and 2) the names of people remaining in $I_{fs}$. Using the same strategy as above, we get $F_{s}$’ name. We extract the full name list from the original text $T_{s}$ using BERT  [devlin2018bert], and then replace $F_{s}$ with inserted person’s name to form the final name list $N_{A}$. Finally, we input $I_{fs}$ and $N_{A}$ into Qwen2-VL to generate aligned text.
$\triangleright$ Face Attribute. Face emotion editing is also considered in our dataset. Our pipeline uses StyleCLIP  [patashnik2021styleclip] and HFGI  [wang2022high_HFGI] for attribute manipulations. Firstly, we analyze facial expressions using an emotion detector  [Arriaga2017RealtimeCN] to determine positive or negative emotions. We then randomly select a method to manipulate the primary face $F_{s}$’ attributes inversely to the classification outcome, producing $I_{fa}$. To ensure diversity, we control manipulation intensity with variable prompts and introduce age modifications. The distribution of face attribute prompts is shown in Fig. 2(b), with $y_{box}$ stored as annotation. Since the characters in $I_{fa}$ have not changed, the paired text in this category is still authentic.
@@
$\triangleright$ Text Fabrication. For the TF category, we also use BERT  [devlin2018bert] to extract the name list $N_{A}$ from the original text. Then, we input $N_{A}$ and the original image into the MLLM to generate narratives that match the implied meaning but are still fabricated.
@@ 3.3 Dataset Statistics
With the above steps, we finally harvest our MDSM dataset $\mathcal{D}_{M}$, a large-scale, 100% semantic-aligned multi-modal benchmark with high-fidelity texts from MLLM. The distribution of manipulation categories is well balanced and consistent with previous datasets, ensuring fair evaluation across manipulation modes (Fig. 2a). Compared with the existing manipulation detection benchmarks in Tab. 1, MDSM has the following advantages: 1)Risk Consideration of MLLM. MDSM acknowledges the emerging challenges posed by MLLMs and utilizes multi-modal methods to create semantically coherent and contextually plausible narratives for manipulated images. This scenario, though underexplored, is a highly significant and timely problem in the modern large model era. 2)Semantic Alignment. MDSM is an aligned multi-modal media manipulation benchmark, which is a significant and more practical scenario for multi-modal manipulation detection. 3)Large Scale. Our MDSM comprises 441,423 samples and is the largest benchmark for detecting and grounding multi-modal manipulation. 4)Diverse Multi-media Sources. The multi-modal media of MDSM sources from diverse media sites, including The Guardian, The New York Times, The Washington Post, USA Today, and the BBC. Consequently, the generality of methods can be assessed via cross-domain evaluation.
@@ 4 Methodology
Fig. 3 illustrates the architecture of our Artifact-aware Manipulation Diagnosis framework (AMD). Built upon Florence-2  [Florence2_CVPR] to leverage real-world knowledge, AMD follows a sequence-to-sequence framework for joint textual detection and grounding. Multimodal inputs are processed through three stages, and outputs localized predictions with textual explanations.
@@
Prompt construction. To adapt the MLLM for the MDSM task while preserving its inherent knowledge, we develop heuristic question(human)-answer(assistant) prompts where the image-question pair serves as input and the text response constitutes AMD’s output:
@@
In this prompt paradigm [^1] Details and examples are given in supplementary materials.,
  • $\textit{Task}>$: Specifies the manipulation detection objective and pairs the input image-text.
  • $\textit{Options}>$: Lists all candidate answers for MDSM task.
  • $\textit{Grounding}>$: Conditionally triggers region localization (via coordinates in brackets) only when image manipulation is detected.
  • $\textit{Response}>$: Encapsulates the correct answers.
  • $\textit{Coordinates}>]$: Optionally encloses tampered region coordinates if the image is altered.

### v4

@@ Abstract
The detection and grounding of multimedia manipulation has emerged as a critical challenge in combating AI-generated disinformation. While existing methods have made progress in recent years, we identify two fundamental limitations in current approaches: (1) Underestimation of MLLM-driven deception risk: prevailing techniques primarily address rule-based text manipulations, yet fail to account for sophisticated misinformation synthesized by multimodal large language models (MLLMs) that can dynamically generate semantically coherent, contextually plausible yet deceptive narratives conditioned on manipulated images; (2) Unrealistic misalignment artifacts: currently focused scenarios rely on artificially misaligned content that lacks semantic coherence, rendering them easily detectable. To address these gaps holistically, we propose a new adversarial pipeline that leverages MLLMs to generate high-risk disinformation. Our approach begins with constructing the MLLM-Driven Synthetic Multimodal (MDSM) dataset, where images are first altered using state-of-the-art editing techniques and then paired with MLLM-generated deceptive texts that maintain semantic consistency with the visual manipulations. Building upon this foundation, we present the Artifact-aware Manipulation Diagnosis via MLLM (AMD) framework featuring two key innovations: Artifact Pre-perception Encoding strategy and Manipulation-Oriented Reasoning, to tame MLLMs for the MDSM problem. Comprehensive experiments validate our framework’s superior generalization capabilities as a unified architecture for detecting MLLM-powered multimodal deceptions. In cross-domain testing on the MDSM dataset, AMD achieves the best average performance, with 88.18 ACC, 60.25 mAP, and 61.02 mIoU scores. The code is available at https://github.com/YcZhangSing/AMD.
Figure 1: Comparison between widely-used $\text{DGM}^{4}$ [41] and MDSM (ours). (left) $\text{DGM}^{4}$ typically treats visual context manipulation as two independent procedures as rule-based text editing and image editing. The absence of context integration often results in poorly aligned samples, which can be readily perceived by the public. (right) In a real-world scenario, we usually face well-wrapped content on purpose. To mimic such a situation, we propose a new manipulation paradigm, which explicitly provides the modified image as well as meta info of facial editing to Multimodal Large Language Model (MLLM). Then we harness MLLM to generate contextually consistent, deceptive texts to form the challenging image-text pairs.
@@ 1 Introduction
Datasets Samples Modality Tasks Training Semantic MLLM
LIAR  [53] 13K
DFIM-HQ  [27] 140K
MEIR  [38] 139k
MiRAGeNews  [15] 15k
COSMOS  [45] 453k
MMFakeBench  [24] 11k
$\text{DGM}^{4}$  [40] 230k
MDSM (Ours) 441k
@@
Generative techniques have become a transformative force in artificial intelligence [54, 33, 9, 12, 5, 58], showing remarkable adaptability across various domains and modalities. These advancements, while enriching multimedia content, also pose significant challenges to information security. In the media industry in particular, maliciously fake content manipulated by such models can profoundly mislead audiences [57, 39, 48]. The unchecked spread of fake media has already negatively affected political, financial, and other sectors [7, 34, 37], gradually becoming a major social issue in current epoch [29, 49, 43].
While various misinformation scenarios have been explored including COSMOS [45], which investigates out-of-context social image-text pairs, and $\text{DGM}^{4}$ [40], which focuses on detecting randomly tampered regions or words. MMFakeBench [24] has made preliminary efforts in detecting semantically aligned fake news. It considers fake images synthesized by text-to-image models to maintain semantic consistency with the rumor texts. However, such semantically aligned instances constitute only 30% of its dataset, and the overall dataset size is merely 11k, which is insufficient to support the training of a robust detection model.
@@
The above analysis reveals two critical limitations in existing research: 1) Neglect of emerging risks from modern Multi-modal Large Language Models (MLLMs): Current paradigms predominantly address rule-based text manipulation, overlooking the sophisticated linguistic capabilities of modern MLLMs. MLLM-generated text exhibits superior fluency and contextual coherence, significantly increasing deception potential and public susceptibility. 2) Semantic misaligned artifacts. Most methods independently manipulate visual and textual elements, producing semantically discordant multimedia. This misalignment not only renders manufactured disinformation too simplistic to effectively deceive the public, but also fails to replicate real-world adversarial behavior, as sophisticated attackers typically maintain meticulous visual-textual consistency to maximize manipulative impact. Both limitations render the multi-modal disinformation scenarios considered in existing works insufficiently realistic.
@@
To address these weaknesses, we take the risk of MLLM-driven fabrication into consideration and focus on detecting the semantic-aligned manipulation. We begin by constructing the MLLM-Driven Synthetic Multi-modal (MDSM) dataset, where both images and text are manipulated in a coordinated manner, resulting in semantically aligned image-text pairs. For image manipulation, we consider the typical Face Swap and Face Attribute editing. For text, We innovatively guide MLLM to generate modality-aligned yet misleading fake narratives using image editing metadata. As shown in Fig. 1(b), after replacing Donald Trump’s face with Micheál Martin’s, we use the swapped name, Micheál Martin, to guide MLLM in generating text, ensuring that the named entity in the text aligns with the image. Following this strategy, we construct over 441k sample pairs.
@@
The alignment of modalities and the authentic texts from MLLMs pose significant challenges for the detection of manipulated media. First, the strategy of perceiving inconsistencies between images and text through contrastive learning, as employed by prior works [59, 40, 51, 42], is ineffective in MDSM where images and text are well-matched already. Merely observing aligned image-text pairs is inadequate for reliable detection. Consequently, external clues and contextual knowledge are essential. Second, existing architectures like ASAP [59] and HAMMER [40], which feature multiple detection and grounding heads, are complex and lack generalizability to unseen media sources. To address these challenges, we propose Artifact-aware Manipulation Diagnosis via MLLM (AMD), which leverages MLLMs’ comprehensive understanding of real-world multimedia and their ability to provide unified textual outputs. And AMD generates detection and grounding results in a coherent, text-based format, offering a more intuitive and generalized solution. In summary, our contributions of this paper are as follows:
@@
  • We propose a strong baseline, i.e., Artifact-aware Manipulation Diagnosis framework (AMD), for the MDSM problem. AMD synergizes artifact pre-perception encoding and manipulation-oriented reasoning to effectively adapt MLLMs for precise manipulation analysis.
@@ 2.1 Multi-Modal Media Source Collection
We use the GoodNews  [6], VisualNews  [13], and N24News  [61] datasets as the Source News Pool $\mathcal{D}_{O}$, which consists of over 2.1M image-text pairs sourced from various real-world news outlets. Given the significant influence of human-centric news among various forms of multi-modal media, we focus on human-centric data for MDSM. $\mathcal{D}_{O}$ is firstly filtered by detecting faces in images with Dlib  [19] and identifying person names in texts with BERT  [10]. Only pairs, $p_{s}=(I_{s},T_{s})$, containing both faces and named entities are used for manipulation. Additionally, we collect the Celebrity Head-shot Dataset $\mathcal{D}_{C}$, which contains about 30k pairs of head-shot images and corresponding names to facilitate the aligned manipulation for Face Swap.
@@ 2.2 Multi-Modal Media Manipulation
In the image modality, two main attacks, Face Swap (FS) and Face Attribute (FA), are employed. For the text modality, we utilize advanced MLLM to generate semantic-aligned Text Fabrication (TF) for the images. Combining these image and text manipulations, we construct a total of five types of multimodal forgeries: FS, FS&TF, FA, FA&TF, and TF. The data generation pipeline is illustrated in Fig. 2, and the detailed generation process is described as follows:
@@
$\triangleright$ Face Swap. Face swap is a critical tool for attackers to forge images of public figures and politicians, posing threats to societal security. We use two representative face swap methods, SimSwap  [8] and e4s  [25], to perform such manipulations. We prioritize modifying larger faces to target the primary subject in the image (Fig. 3c). Given a source image $I_{s}$, we randomly choose one of the two methods and replace the largest face $F_{s}$ in $I_{s}$ with a face $F_{c}$ from $\mathcal{D}_{C}$, generating a manipulated face swap sample $I_{fs}$. The bounding box $y_{box}=\{x_{1},y_{1},x_{2},y_{2}\}$ of the swapped face and the name of $F_{c}$ are recorded. To keep the image-text aligned, corresponding processing is also done on the authentic text. We use MLLM to identify the name of $F_{s}$, as shown in Fig. 2, $F_{s}$ is identified as Trump. We then refine the authentic text for FS category by replacing this name with $F_{c}$’s name.
$\triangleright$ Face Swap and Text Fabrication. We use Qwen2-VL  [50] to generate consistent but misleading narratives. This requires knowing: 1) the inserted person’s name, and 2) the names of people remaining in $I_{fs}$. Using the same strategy as above, we get $F_{s}$’ name. We extract the full name list from the original text $T_{s}$ using BERT  [10], and then replace $F_{s}$ with inserted person’s name to form the final name list $N_{A}$. Finally, we input $I_{fs}$ and $N_{A}$ into Qwen2-VL to generate aligned text.
$\triangleright$ Face Attribute. Face emotion editing is also considered in our dataset. Our pipeline uses StyleCLIP  [32] and HFGI  [52] for attribute manipulations. Firstly, we analyze facial expressions using an emotion detector  [28] to determine positive or negative emotions. We then randomly select a method to manipulate the primary face $F_{s}$’ attributes inversely to the classification outcome, producing $I_{fa}$. To ensure diversity, we control manipulation intensity with variable prompts and introduce age modifications. The distribution of face attribute prompts is shown in Fig. 3(b), with $y_{box}$ stored as annotation. Since the characters in $I_{fa}$ have not changed, the paired text in this category is still authentic.
@@
$\triangleright$ Text Fabrication. For the TF category, we also use BERT  [10] to extract the name list $N_{A}$ from the original text. Then, we input $N_{A}$ and the original image into the MLLM to generate narratives that match the implied meaning but are still fabricated.
@@ 2.3 Dataset Statistics
With the above steps, we finally harvest our MDSM dataset $\mathcal{D}_{M}$, a large-scale, high semantic-aligned multi-modal benchmark with high-fidelity texts from MLLM. The distribution of manipulation categories is well balanced and consistent with previous datasets, ensuring fair evaluation across manipulation modes (Fig. 3a). Compared with the existing manipulation detection benchmarks in Tab. 1, MDSM has the following advantages: 1)Risk Consideration of MLLM. MDSM acknowledges the emerging challenges posed by MLLMs and utilizes multi-modal methods to create semantically coherent and contextually plausible narratives for manipulated images. This scenario, though underexplored, is a highly significant and timely problem in the modern large model era. 2)Semantic Alignment. MDSM is a high aligned multi-modal media manipulation benchmark, which is a significant and more practical scenario for multi-modal manipulation detection. 3)Large Scale. Our MDSM comprises 441,423 samples and is the largest benchmark for detecting and grounding multi-modal manipulation. 4)Diverse Multi-media Sources. The multi-modal media of MDSM sources from diverse media sites, including The Guardian, The New York Times, The Washington Post, USA Today, and the BBC. Consequently, the generality of methods can be assessed via cross-domain evaluation.
@@ 3 Methodology
Fig. 4 illustrates the architecture of our Artifact-aware Manipulation Diagnosis framework (AMD). Built upon Florence-2  [5] to leverage real-world knowledge, AMD follows a sequence-to-sequence framework for joint textual detection and grounding. Multimodal inputs are processed through three stages, and outputs localized predictions with textual explanations.
@@
Prompt construction. To adapt the MLLM for the MDSM task while preserving its inherent knowledge, we develop heuristic question(human)-answer(assistant) prompts:
@@
[…]